### v1

# Search for Dark Matter Annihilation to gamma-rays from SPT-SZ selected Galaxy Clusters

Abstract
We search for dark matter annihilation from galaxy clusters in the energy range from 1-300 GeV using nearly 16 years of Fermi-LAT data. For this purpose, we use 350 galaxy clusters selected from the 2500 SPT-SZ survey. We model the dark matter distribution using the NFW profile for the main halo along with the Einasto profile for the substructure. The largest signal is seen for the cluster SPT-CL J2021-5257 with a significance of around . The best-fit dark matter mass and annihilation cross-section for this cluster are equal to GeV and for the annihilation channel. However, this central estimate is in conflict with the limits on annihilation cross-section from dwarf spheroidal galaxies and hence cannot be attributed to dark matter annihilation. Three other clusters show significance between , whereas all the remaining clusters show null results. The most stringent 95% c.l. upper limit, which we obtain among the clusters with significance is from SPT-CL J2300-5331, given by for a dark matter mass of 10 GeV corresponding to the annihilation channel.
I Introduction
The dark matter problem is one of the most outstanding unsolved problems in Astrophysics Jungman et al. (1996); Bertone et al. (2005); Bertone and Hooper (2018); Bosma (2023). Although a large number of observations have pointed to a concordance model of the universe, which consists of about 25% cold dark matter Planck Collaboration et al. (2020), the identity of the cold dark matter candidate is unknown. One of the most well motivated and widely studied dark matter candidate is a weakly interacting massive particle (WIMP), which is a hypothetical stable particle, assumed to be a Majorana fermion with masses in the GeV-TeV range and weak-scale interactions Lee and Weinberg (1977); Jungman et al. (1996); Bertone et al. (2005). The most well motivated WIMP candidate is the lightest supersymmetric particle, which is usually the neutralino Ellis et al. (1984). Such a hypothetical particle has velocity averaged cross-section given by . This cross-section yields the correct relic abundance, and enables the WIMP to decouple from the rest of the universe at non-relativistic velocities, thereby providing all the right properties needed for a cold dark matter candidate Jungman et al. (1996); Bertone et al. (2005); Steigman et al. (2012). There are two kinds of experiments designed to look for signatures of WIMPs. The first type of search involves direct detection of WIMPs, where one looks for signatures of WIMP scattering with nuclei in an underground experiment Schumann (2019). The second search is referred to as indirect detection Cirelli (2012); Donato (2014), where one looks for signatures of WIMP annihilation into secondary particles such as such as neutrinos Desai et al. (2004), positrons Cholis et al. (2009), anti-protons Heisig (2021), radio waves Chan (2021), and gamma-rays Funk (2015).
In this work, we search for dark matter annihilation to gamma-rays using galaxy clusters. Galaxy clusters are the most massive gravitationally bound and virialized structures in the Universe and act as a unique laboratory to probe cosmology Kravtsov and Borgani (2012); Allen et al. (2011); Vikhlinin et al. (2014) and fundamental Physics Böhringer and Chon (2016); Desai (2018); Bora and Desai (2021a, b); Bora et al. (2022). Most of the mass of galaxy clusters (80-85%) is made up of dark matter Allen et al. (2011). Therefore, galaxy clusters are the largest reservoirs of dark matter and hence provide excellent targets in looking for signatures of dark matter annihilation Colafrancesco et al. (2006).
The Fermi Gamma-ray Space Telescope which was launched in 2008 is currently the most sensitive telescope for gamma-ray astronomy. The Large Area Telescope (LAT) is one of the two instruments onboard this detector. Fermi-LAT is sensitive to high energy gamma rays from various astrophysical sources. It is a pair-conversion telescope that is sensitive to photons between the energy range of 20 MeV to more than 300 GeV Atwood et al. (2009) and is sensitive to dark matter annihilation if the dark matter is a WIMP with masses in the GeV range, since the photon energy from the annihilation products should be within the detector sensitivity.
A large number of searches have been done looking for dark matter annihilation in galaxy clusters using Fermi-LAT Ackermann et al. (2010); Ando and Nagai (2012); Huang et al. (2012); Han et al. (2012); Hektor et al. (2013); Anderson et al. (2016); Liang et al. (2016); Ackermann et al. (2015); Lisanti et al. (2018); Thorpe-Morgan et al. (2021); Shen et al. (2021); Di Mauro et al. (2023); Song et al. (2024). We briefly summarize these works. The very first search for dark matter annihilation from clusters was done using 11 months of Fermi-LAT data from six X-ray selected clusters in the HIFLUGCS catalog Ackermann et al. (2010). No significant gamma-ray detection was reported in this analysis. Subsequently, null results were reported based on a search from 49 galaxy clusters located at high galactic latitude using 2.8 years of Fermi-LAT data Ando and Nagai (2012). The most stringent upper limit from this analysis was from the Fornax cluster, given by Ando and Nagai (2012). Then, another search for dark matter annihilation and decay was done from eight nearby galaxy clusters (both individually and from a stacked analysis) using three years of Fermi-LAT data Huang et al. (2012). No evidence for gamma-ray emission at significance due to dark matter decay or annihilation was found from this analysis. Limits on the annihilation cross-section of around were then obtained Huang et al. (2012). A similar search for dark matter annihilation from three nearby clusters, namely, VIRGO, Coma, Fornax was done using 45 months of Fermi-LAT data Han et al. (2012). After removing the galactic and extragalactic gamma-ray foregrounds, a residual diffuse emission was found, which was attributed to additional point sources not included in the Fermi-LAT two-year catalog Han et al. (2012). After accounting for these, no significant extended emission was seen from the aforementioned three clusters and upper limits on the cross section for dark matter annihilation from the three clusters were obtained, which were more stringent than those obtained from Milky way dwarf galaxies Han et al. (2012). Subsequently, Hektor et al. (2013) found evidence for a double-peaked line at 110 and 130 GeV from 18 nearby brightest galaxy clusters with more than significance using 218 weeks of Fermi-LAT data, which they attributed to dark matter annihilation. However, this signal could not be confirmed in a subsequent follow-up analysis using enhanced exposure Anderson et al. (2016); Liang et al. (2016). A dedicated search for dark matter annihilation and point source emission from the VIRGO cluster was undertaken by the Fermi-LAT collaboration using three years of data, which reported null results Ackermann et al. (2015). This work then calculated limits on velocity averaged DM annihilation cross-section. Then in 2016, two independent groups did a search around the same time from the same set of 16 galaxy clusters from the HIFLUGCS sample with the largest -factors, using five and seven years of data respectively Anderson et al. (2016); Liang et al. (2016). The first analysis Anderson et al. (2016) did not find any significant excess in the energy range between 10 to 400 GeV and upper limits on the velocity averaged annihilation cross-section of around were obtained. The second analysis found a line at 43 GeV after stacking with a significance of Liang et al. (2016). However, since the backgrounds were not fully understood, they did not claim a detection and set a conservative limit on the the velocity averaged annihilation cross-section Liang et al. (2016). A search for dark matter annihilation from galaxy groups (lower mass analogs of galaxy clusters) having large -factors using 413 weeks of Fermi-LAT data Lisanti et al. (2018). No evidence for dark matter annihilations was found and thermal relic cross sections for dark matter masses below approximately 30 GeV to annihilation channel were excluded at 95% c.l. Then, another search for dark matter annihilation using 12 years of Fermi-LAT data from five nearby clusters (Centaurus, Coma, VIRGO, Fornax, and Perseus) reported null results and constraints on annihilation cross-section into three different channels were reported, which are given between , depending on the WIMP mass Thorpe-Morgan et al. (2021). A search for line and box-shaped signals was done in Shen et al. (2021) using 7.1 years of Fermi-LAT by stacking the same set of 16 clusters analyzed in Anderson et al. (2016); Liang et al. (2016). This work also found the 42 GeV line first reported in Liang et al. (2016), although its significance was reduced. Consequently , 95% c.l. upper limits on the thermally averaged DM annihilation cross-section to photons between were set Shen et al. (2021). About a year ago, Di Mauro et al. (2023) (D23, hereafter) used 12 years of Fermi-LAT data to look for dark matter annihilation and decay from 49 clusters selected from the HIFLUGCS sample. The stacking analysis revealed a signal at 2.5-3.0 significance. However, the best-fit values of the mass and cross-section are in tension with the null results obtained using dark matter searches from dwarf spheroidal galaxies. Therefore, it was concluded that the signal is most likely due to cosmic ray collisions with the gas and photon fields within the cluster Di Mauro et al. (2023). Most recently, a constraint on the lifetime of very heavy dark matter with masses between and GeV was set using 14 years of Fermi-LAT data, based on the analysis of seven clusters Song et al. (2024).
Complementary to the aforementioned target searches from individual clusters or a stacked analysis, a cross-correlation analysis between the 9-year Fermi-LAT diffuse gamma-ray map and four different galaxy cluster catalogs was done to look for dark matter annihilations, which reported null results Tan et al. (2020). In additional to galaxy clusters, a large number of works have also looked for dark matter annihilations and decay from Galactic center, dwarf spheriodal galaxies as well as Milky way satellites.
In this work, we look for WIMP dark matter annihilation from the galaxy clusters detected by the South Pole Telescope (SPT) 2500 sq. deg survey, which have been detected using the Sunyaev-Zeldovich (SZ) effect Sunyaev and Zeldovich (1972), as a followup to our previous work which looked for point-source gamma-ray emission using 15 years of Fermi-LAT data Manna and Desai (2024a). We mainly follow the same methodology as D23. In our analysis, we have assumed a flat CDM cosmology with and . This manuscript is structured as follows: In Section II, we define the cluster catalog utilized in our manuscript. Section III provides a comprehensive description of the halo and subhalo modelling. Subsequently, in Section IV, we detail the analysis process employed in our study. Finally, in Section V, we present our results and present our conclusions in Section VI.
II SPT-SZ cluster catalog
For our analysis, we use the SZ selected cluster catalog detected by SPT, which provides a mass-limited sample. The SPT is a 10-meter telescope located at the South Pole that has imaged the sky at three different frequencies: 95 GHz, 150 GHz, and 220 GHz with an angular resolution of about 1 arcminute Carlstrom et al. (2011). SPT completed a 2500 square-degree survey between 2007 and 2011 to detect galaxy clusters using the SZ effect. This 2500 sq. degree SPT-SZ survey detected 677 confirmed galaxy clusters with SNR greater than 4.5, corresponding to a mass threshold of up to redshift of 1.8 Bleem et al. (2015a); Bocquet et al. (2019)111https://pole.uchicago.edu/public/data/sptsz-clusters/2500d_cluster_sample_Bocquet19.fits. The SPT cluster redshifts were obtained using a dedicated optical and infrared follow-up campaign from pointed imaging and spectroscopic observations Song et al. (2012); Ruel et al. (2014) in conjunction with data from optical imaging surveys such as the Blanco Cosmology Survey Desai et al. (2012) and Dark Energy Survey Saro et al. (2015). We have previously done a search for gamma-rays induced due to astrophysical processes from 300 clusters from this sample, ranked according to Manna and Desai (2024a). Here, denotes the total mass enclosed within a sphere with a mean density of 500 times the critical density of the universe at the cluster’s redshift, and represents the cluster’s redshift Bocquet et al. (2019). In this work, we conduct a search for gamma-ray emission from dark matter annihilation for a sample of 350 galaxy clusters from the aforementioned catalog, the clusters are ranked in decreasing order of their mass ratio, .
III Dark Matter Analysis
The expected gamma-ray flux from the annihilation of WIMPs can be computed using the following equation Di Mauro et al. (2023); Charbonnier et al. (2012); Bonnivard et al. (2016); Hütten et al. (2019); Cirelli et al. (2011)222We follow the same notation as D23:
Here, is the WIMP dark matter induced differential gamma-ray flux per unit energy and solid angle as a function of energy (), solid angle (), and line of sight (l.o.s.). The term is the gamma-ray energy spectrum per annihilation as a function of energy and represents the particle physics component of the gamma-ray flux, which encapsulates the spectral characteristics of the WIMP annihilation such as the properties of the dark matter particles and the annihilation channels involved. The differential gamma-ray flux is then computed as follows Di Mauro et al. (2023):
Here, denotes the WIMP photon spectrum, whose calculation is discussed in Cirelli et al. (2011). The mass of the WIMP, denoted as , determines the energy scale of the gamma rays produced during annihilation. The thermally-averaged annihilation cross-section, represented by , quantifies the likelihood of dark matter particles annihilating when they encounter each other. We can define the astrophysical -factor as the line-of-sight integral of the squared DM density profile over the solid angle at a given coordinate in the l.o.s. directions Pace and Strigari (2019); Di Mauro et al. (2023). The -Factor can be mathematically written as:
Here, is the solid angle over which the integration is performed, is the total dark matter density profile as a function of the distance from the centre of the system, and the integration is carried out along the line-of-sight (l.o.s.). The solid angle is related to the integration angle as:
where is the angle between the line-of-sight and the direction pointing toward the centre of the astrophysical system. In our study, the integration angle is considered as 0.1∘. In the context of modelling the dark matter distribution within galaxy clusters, a common approach is to consider the total dark matter density profile as the sum of two components Di Mauro et al. (2023):
The first term on the right-hand side, , represents the smooth, large-scale distribution of dark matter within the main halo of the galaxy clusters, whereas the second term () accounts for the contribution of the population of subhalos within the main halo of the galaxy cluster, according to the standard CDM cosmological model Prada et al. (2012); Klypin et al. (2011); Gao et al. (2004). Subhalos are smaller-scale dark matter structures that are gravitationally bound within the larger main halo Kuhlen et al. (2012); Zavala and Frenk (2019). Since individual subhalos cannot be resolved, we use a statistical description of the subhalos to calculate the annihilation signal.
In the next subsection, we discuss the modelling of the smooth halo and subhalo dark matter component. This modelling process allows us to calculate the expected fluxes arising from DM annihilation within these clusters.
III.1 Modelling the Main Halo
For our work, we modelled the main smooth dark matter halo using the Navarro-Frenk-White (NFW) density profile Navarro et al. (1996, 1997); Zhao (1996), which can be written as follows:
Here, is the characteristic dark matter density for the NFW profile and is the NFW scale radius. The NFW profile is usually reparameterized in terms of the halo mass and concentrations as described below.
The first step in obtaining the halo concentration involves choosing a mass proxy. For this purpose, we use , which is the cluster mass at a spherical over-density with a density 200 times the critical density of the Universe. We have obtained values from the SPT-SZ 2500 square-degree catalogue Bocquet et al. (2019); Bleem et al. (2015b). In accordance with Di Mauro et al. (2023); White (2001), we define the virial radius, which can be estimated from as follows:
where is the overdensity factor, which is 200 in this case. The critical density, , is calculated using the Hubble parameter , which is the rate at which the universe expands and is given by:
where for flat CDM. We now introduce the halo concentration () defined as .
Next, we need to determine for every cluster. Although, an extensive weak lensing campaign has been undertaken for the SPT clusters for precise mass measurements and cosmological analyses Bocquet et al. (2019) (and references therein), no concentration measurements are available on a per cluster basis for the SPT sample. Therefore, we use the concentration-mass () relations from cosmological simulations for modelling the dark matter halo. Similar to D23, we use the concentration-mass relations from Sánchez-Conde and Prada (2014), which have been shown to match the observation results across a wide range of halo masses, from dwarf spheroidal galaxies to galaxy clusters. This mass-concentration relations allows us to obtain from the estimated for every SPT cluster.
III.2 Modelling the subhalo
Structure formation follows a bottom-up hierarchical process in the standard CDM cosmological model Frenk and White (2012). This means that the smallest structures, known as subhalos, form first. Over time, these subhalos merge and accrete matter, leading to the formation of larger and larger structures, such as galaxy groups and clusters Zavala and Frenk (2019). Galaxy clusters are expected to host a large number of subhalos.
We shall now discuss the modelling of the subhalo population along the same lines as D23. The subhalos account for the smaller-scale clumps and inhomogeneities within the halo.
For this purpose, we use the CLUMPY v3 software Charbonnier et al. (2012); Bonnivard et al. (2016); Hütten et al. (2018, 2019)333https://clumpy.gitlab.io/CLUMPY/v3.1.1/_downloads/975332ce6631f0956830ed27431f3b25/CLUMPY_v2018.06.CPC.tar.gz where the Einasto profile Einasto (1965) is used to describe the density distribution of subhalos.
The distribution of subhalos within a main halo is influenced by both their distance from the host center and also their mass. The total distribution of subhalos can be expressed as the product of three uncorrelated probability distribution functions (PDFs) and a normalization factor Di Mauro et al. (2023); Charbonnier et al. (2012); Bonnivard et al. (2016); Hütten et al. (2018) as follows:
In the above equation, refers to the expected number of subhalos within the virial radius of main halo, and with is the probability distribution in each of the respective domains, normalized to 1. Here, corresponds to the volume of the main halo, pertains to the distribution of subhalo masses, and represents the subhalo concentration. Although, numerical cosmological simulations have significantly advanced our understanding of halo substructures Zavala and Frenk (2019), several questions still remain unresolved, such as the minimum mass for clump formation Colafrancesco et al. (2006), the effects of tidal stripping on subhalo survival Aguirre-Santaella et al. (2023), and the exact shape of subhalo DM density profiles Errani and Navarro (2021). These uncertainties impact the calculation of the DM-induced -ray flux. In this work, we follow the same prescription as D23 to account for each of the terms in Eq. 11, which we describe below:
Spatial distribution PDF (): Since the main halo is spherically symmetrical, the distribution of subhalos within it depends only on their distance from the central point of the host halo. It describes the probability of finding a subhalo at a given distance from the host halo centre and provides the spatial distribution of substructures in a host halo, which in our study are the galaxy clusters. We modelled it using the Einasto profile.
Mass distribution PDF (): Subhalo mass function can be defined as:
(12) We have used = 1.9 in our study, similar to Di Mauro et al. (2023); Springel et al. (2008). It represents a more conservative approach as it implies lesser number of subhalos. It also aligns with the results reported in Zavala and Frenk (2019).
Concentration distribution PDF (): As reported in Aguirre-Santaella et al. (2023); Springel et al. (2008); Errani and Peñarrubia (2020); Errani and Navarro (2021); Kazantzidis et al. (2004); Peñarrubia et al. (2010), subhalos experience tidal forces that generally result in significant mass loss, particularly in their outer regions. As a result, subhalos tend to be more concentrated compared to the main halos with equivalent mass Moliné et al. (2017, 2023); Newton et al. (2022). We have used () relation defined in Moliné et al. (2017) for the subhalos, similar to D23. It accounts for the spatial dependence of the subhalos within the main halos.
In order to account for the uncertainties in the above factors, D23 considered three benchmark models. In this work, we choose the parameters corresponding to the MED model in D23. The smallest subhalo mass is assumed to be and the largest subhalo mass is 0.01% of the halo mass. Finally, we specify the number of multilevel substructures to be two, as in D23. All the relevant properties used for modelling of dark matter halos and subhalos are summarized in Table 1. The distribution of -factors for SPT-SZ clusters using this model for the dark matter halo and substructure can be found in Figure 1. The values for individual clusters are listed in Table LABEL:tab:TableII.
IV Fermi-LAT Analysis
IV.1 Data selection
We conducted our search using 15.7 years of Fermi-LAT data from August 5, 2008 to April 1, 2024, (MET 239587201-733622405). For each cluster, a circular region of interest (ROI) with a 10∘ radius centered on the cluster was defined. The analysis was restricted to events within this ROI with energies ranging from 1 GeV to 300 GeV with 37 logarithmic energy bins per decade. We excluded lower energy events (below 1 GeV) due to the degraded Point Spread Function (PSF) at those energies Ackermann et al. (2012a, b). The binned maximum-likelihood method was implemented using the P8R3_ULTRACLEANVETO_V3 instrument response functions (IRFs) and Pass 8 ULTRACLEANVETO ("FRONT+BACK") class events Atwood et al. (2013).
We also applied the filters, DATA_QUAL > 0 and LAT_CONFIG == 1, which were utilized to eliminate low-quality data. The rock angle was constrained to abs(rock angle) and |b| to mitigate particle contamination. To reduce atmospheric interference, a zenith angle cut of 90∘ was imposed on the events. We also removed all clusters within of the galactic plane. The spatial binning of the data was performed at a resolution of 0.2 pixels, which helps to reduce noise and improve the signal-to-noise ratio. Our background model incorporated all the gamma-ray sources from the fourth Fermi-LAT catalogue (4FGL-DR4) Ballet et al. (2023), encompassing both point-like and extended sources. To account for the pervasive, low-intensity gamma-ray emission from the Milky Way galaxy itself (diffuse emission), we employed the Galactic diffuse emission model (gll_iem_v07.fits). Additionally, an isotropic component (iso_P8R3_ULTRACLEANVETO_V3_v1.txt) is included to represent the isotropic gamma-ray background from various unresolved extragalactic sources. The analysis pipeline employed the Fermitools (v2.2.0) software within the Fermi Science Support Center’s Fermibottle Docker environment Wood et al. (2017).
IV.2 Analysis Method
Our analysis followed the standard binned-likelihood method developed by the Fermi-LAT collaboration for likelihood calculations and model fitting Atwood et al. (2009). We incorporated sources from the 4FGL catalog up to beyond the defined region of interest into our model, freezing all their parameters to the catalog values. This was done to minimize bias from the potential presence of bright sources outside the targeted area and to account for the LAT’s poor point spread function (PSF) at low energies around 0.1 GeV Thorpe-Morgan et al. (2021); Manna and Desai (2024a). In the first phase of the analysis, the spectral parameters for all the free sources were determined. During the second phase, we then fixed the spectral parameters of all sources, except for the normalizations to their best-fit values. Then, we included a template for potential dark matter annihilation emission in the model which we modeled as a diffuse source with a spatial distribution proportional to the -factor calculated using CLUMPY and specified in Table LABEL:tab:TableII. The spectral component of the signal was given by the spectrum of annihilating dark matter for the specified channel utilizing the DMFitFunction Jeltema and Profumo (2008) within Fermitools similar to Thorpe-Morgan et al. (2021).
We used the gtapps tool for our analysis and employed the Maximum Likelihood Estimation (MLE) technique to identify the model parameters that best match the source’s spectrum and location. The gtlike tool performs a binned likelihood analysis on Fermi-LAT gamma-ray data. It achieves this by comparing a model of the gamma-ray sky with the actual observations and calculating the likelihood that the model explains the data. To identify gamma-ray sources and assess their significance, we employ the Test Statistic (TS) calculated using the gttsmap tool. This statistic, defined as Mattox et al. (1996)
compares the likelihoods of two scenarios: where is a model without the source (null hypothesis) and is a model including the source at a specific location (alternative hypothesis). Wilks’ theorem tells us that for high photon counts, the TS for the null hypothesis behaves similarly to a chi-squared distribution Wilks (1938). The detection significance is thereby nominally obtained by taking the square root of TS. However, D23 has shown using simulations, that for the null hypothesis the TS statistics has larger tails, which are not expected in the distribution. Therefore, the actual detection significance would be expected to be smaller than the square root of TS. Nevertheless, we shall quote the significance as the square root of TS for our results. We note that the same TS statistic is also used in neutrino and soft gamma-ray astrophysics to evaluate the significance of detections Pasumarti and Desai (2022); Manna and Desai (2024b).
Parameter | Value |
Smallest subhalo mass | |
Biggest subhalo mass | 0.01 % of host mass |
Number of multilevel substructures | 2 |
Mass-concentration model for Halos | Sánchez-Conde and Prada (2014) |
Mass-concentration model for subhalos | Moliné et al. (2017) |
profile of subhalo distribution in host | EINASTO |
Slope of power law subhalo mass spectrum | 1.9 |
Fraction of host halo mass bound in subhalos | 0.1 |
Branching Ratio Channel | , |
Spectrum Model | Cirelli et al. (2011) |
Dark Matter Particle | Majorana |
Spatial -Factor Integration angle () |
V Results
We conducted the analysis for the clusters in our sample using and annihilation channels. The TS values for all the 350 galaxy clusters can be found in Table LABEL:tab:TableII. Among all the clusters, SPT-CL J2021-5257 shows the maximum TS value of 9, corresponding to 3 emission using the dark matter template as shown in Figure 2 for the energy range between 1 GeV and 300 GeV. We also performed the analysis for SPT-CL J2021-5257 by extending the lower energy range to 100 MeV and found that TS value decreases to 7, corresponding to 2.6 significance. Other clusters, such as SPT-CL J2012-5649, SPT-CL J0124-5937, and SPT-CL J2300-5331 show TS values corresponding to significance, suggesting a weaker detection significance. The variation of TS values as a function of the dark matter mass for the aforementioned clusters with significance can be found in Figure 3 for annihilation channel . The best fit WIMP mass for SPT-CL J2021-5257 is found to be GeV, whereas the best-fit value of for this cluster assuming a annihilation channel. For SPT-CL J2012-5649, SPT-CLJ 0124-5937, and SPT-CL J2300-5331, the best-fit masses were found to be GeV, GeV, and GeV, respectively. The corresponding best-fit values of for the channel were , and , respectively. The best-fit dark matter mass for SPT-CL J2021-5257 for the annihilation channel is found to be GeV, whereas the best-fit annihilation cross-section is given by . Similarly, for SPT-CL J2012-5649, SPT-CLJ0124-5937, and SPT-CL J2300-5331 we found the best-fit mass as , , and GeV, respectively, whereas the best-fit values of are equal to , and , respectively for this annihilation channel. The variation of TS with mass for the annihilation channel for all the four clusters can be found in Fig. 4.
As pointed out in D23, these values are in conflict with the upper limits obtained from null searches for dark matter annihilation from Milky Way dwarf spheroidal galaxies, whose 95% c.l. limits on the velocity-averaged annihilation cross-section are around Di Mauro and Winkler (2021). Therefore, although our significance does not cross the threshold, these enhanced TS values corresponding to (2-3) significance cannot be by-products of dark matter annihilation. These are either statistical fluctuations or caused by some astrophysical processes.
Therefore, we calculate upper limits for the dark matter annihilation cross-section for these clusters. We show the 95% c.l. upper limits on the annihilation cross-section of WIMP dark matter mass for the annihilation channel in Figure 5 and for the channel in Figure 6. Our upper limits are mostly in agreement with other works, which have obtained limits on annihilation cross-section Di Mauro et al. (2023); Ackermann et al. (2010); Han et al. (2012); Thorpe-Morgan et al. (2021). Among all the clusters with significance , we found the most stringent limit for SPT-CL J2300-5331, viz. for GeV and annihilation channel and for GeV and annihilation channel. We note among the clusters with signifiance is SPT-CL J2012-5649 (TS = 6.2), which is spatially coincident with the merging cluster Abell 3667. We note that this cluster was detected with the highest significance () in our previous work on gamma-ray searches using a point source template Manna and Desai (2024a).
All other clusters have significance less then .
These include the merging clusters in our sample such as Bullet Cluster (SPT-CL J0658-5556) and El Gordo Cluster (SPT-CL J0102-4915) with TS values of 1.5 and 1.9, respectively, corresponding to no significant excess. Since the last 100 clusters in our sample (sorted according to ) showed no significance we restricted our analysis to 350 clusters and did not go beyond that.
VI Conclusions
In this work, we searched for gamma-ray emission between 1 and 300 GeV from dark matter annihilation in galaxy clusters. We analyzed 350 clusters selected from the SPT-SZ 2500 square degree survey, which provides a mass-limited sample. The analysis utilized 15.7 years of data from the Fermi-LAT telescope and employed the DMFIT template. Notably, we identified a emission signal from the galaxy cluster SPT-CL J2021-5257. The best fit value for for SPT-CL J2021-5257 is found to be and best fit dark matter mass as GeV assuming annihilation channel and for annihilation channel we found the best fit value as and best fit dark matter mass as GeV. The TS map for this cluster can be found in Fig. 2. Although prima-facie, we found detection evidence from SPT-CL J2021-5257, the estimated annihilation cross-section is in conflict with upper limits from Milky way dwarf spheroidal galaxies. Therefore, we conclude that the enhanced signal seen for SPT-CL J2021-5257 cannot be due to dark matter annihilation. We obtained a marginal significance of around 2 for three other clusters: SPT-CL J2012-5649, SPT-CL J0124-5937 and SPT-CL J2300-5331. All the remaining clusters showed null results and the TS values were consistent with background.
The most stringent upper limits we obtain for annihilation to is given by SPT-CL J2300-5331 as for GeV and annihilation channel and for GeV and annihilation channel.
Moving forward, we plan to expand this analysis to encompass all galaxy clusters detected in ongoing X-ray and SZ surveys such as eROSITA Kluge et al. (2024) and SPTPol Bleem et al. (2024).
Acknowledgements.
We are grateful to the Fermi-LAT and CLUMPY teams for their invaluable contributions in making the data and analysis codes publicly available. Their expertise and willingness to answer our queries were instrumental in this research. This work was motivated after a stimulating talk by Mattia Di Mauro at IIT Hyderabad. We also acknowledge the National Supercomputing Mission (NSM) for providing essential computing resources on the ‘PARAM SEVA’ supercomputer at IIT Hyderabad. This initiative is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST) of the Government of India. Finally, we thank the Smithsonian Astrophysical Observatory for developing SAOImageDS9, which played a valuable role in our analysis.References
- Jungman et al. (1996) G. Jungman, M. Kamionkowski, and K. Griest, Physics Reports 267, 195 (1996), eprint hep-ph/9506380.
- Bertone et al. (2005) G. Bertone, D. Hooper, and J. Silk, Physics Reports 405, 279 (2005), eprint hep-ph/0404175.
- Bertone and Hooper (2018) G. Bertone and D. Hooper, Reviews of Modern Physics 90, 045002 (2018), eprint 1605.04909.
- Bosma (2023) A. Bosma, arXiv e-prints arXiv:2309.06390 (2023), eprint 2309.06390.
- Planck Collaboration et al. (2020) Planck Collaboration, N. Aghanim, Y. Akrami, M. Ashdown, J. Aumont, C. Baccigalupi, M. Ballardini, A. J. Banday, R. B. Barreiro, N. Bartolo, et al., Astron. & Astrophys. 641, A6 (2020), eprint 1807.06209.
- Lee and Weinberg (1977) B. W. Lee and S. Weinberg, Phys. Rev. Lett. 39, 165 (1977).
- Ellis et al. (1984) J. Ellis, J. S. Hagelin, D. V. Nanopoulos, K. Olive, and M. Srednicki, Nuclear Physics B 238, 453 (1984).
- Steigman et al. (2012) G. Steigman, B. Dasgupta, and J. F. Beacom, Phys. Rev. D 86, 023506 (2012), eprint 1204.3622.
- Schumann (2019) M. Schumann, Journal of Physics G Nuclear Physics 46, 103003 (2019), eprint 1903.03026.
- Cirelli (2012) M. Cirelli, Pramana 79, 1021 (2012), eprint 1202.1454.
- Donato (2014) F. Donato, Physics of the Dark Universe 4, 41 (2014).
- Desai et al. (2004) S. Desai, Y. Ashie, S. Fukuda, Y. Fukuda, K. Ishihara, Y. Itow, Y. Koshio, A. Minamino, M. Miura, S. Moriyama, et al., Phys. Rev. D 70, 083523 (2004), eprint hep-ex/0404025.
- Cholis et al. (2009) I. Cholis, D. P. Finkbeiner, L. Goodenough, and N. Weiner, JCAP 2009, 007 (2009), eprint 0810.5344.
- Heisig (2021) J. Heisig, Modern Physics Letters A 36, 2130003-414 (2021), eprint 2012.03956.
- Chan (2021) M. H. Chan, Galaxies 9, 11 (2021).
- Funk (2015) S. Funk, Proceedings of the National Academy of Science 112, 12264 (2015), eprint 1310.2695.
- Kravtsov and Borgani (2012) A. V. Kravtsov and S. Borgani, Ann. Rev. Astron. Astrophys. 50, 353 (2012), eprint 1205.5556.
- Allen et al. (2011) S. W. Allen, A. E. Evrard, and A. B. Mantz, Ann. Rev. Astron. Astrophys. 49, 409 (2011), eprint 1103.4829.
- Vikhlinin et al. (2014) A. A. Vikhlinin, A. V. Kravtsov, M. L. Markevich, R. A. Sunyaev, and E. M. Churazov, Physics Uspekhi 57, 317-341 (2014).
- Böhringer and Chon (2016) H. Böhringer and G. Chon, Modern Physics Letters A 31, 1640008 (2016), eprint 1610.02855.
- Desai (2018) S. Desai, Physics Letters B 778, 325 (2018), eprint 1708.06502.
- Bora and Desai (2021a) K. Bora and S. Desai, JCAP 2021, 012 (2021a), eprint 2008.10541.
- Bora and Desai (2021b) K. Bora and S. Desai, JCAP 2021, 052 (2021b), eprint 2104.00974.
- Bora et al. (2022) K. Bora, R. F. L. Holanda, S. Desai, and S. H. Pereira, European Physical Journal C 82, 17 (2022), eprint 2106.15805.
- Colafrancesco et al. (2006) S. Colafrancesco, S. Profumo, and P. Ullio, Astron. & Astrophys. 455, 21 (2006), eprint astro-ph/0507575.
- Atwood et al. (2009) W. B. Atwood, A. A. Abdo, M. Ackermann, W. Althouse, B. Anderson, M. Axelsson, L. Baldini, J. Ballet, D. L. Band, G. Barbiellini, et al., Astrophys. J. 697, 1071 (2009), eprint 0902.1089.
- Ackermann et al. (2010) M. Ackermann, M. Ajello, A. Allafort, L. Baldini, J. Ballet, G. Barbiellini, D. Bastieri, K. Bechtol, R. Bellazzini, R. D. Blandford, et al., JCAP 2010, 025 (2010), eprint 1002.2239.
- Ando and Nagai (2012) S. Ando and D. Nagai, JCAP 2012, 017 (2012), eprint 1201.0753.
- Huang et al. (2012) X. Huang, G. Vertongen, and C. Weniger, JCAP 2012, 042 (2012), eprint 1110.1529.
- Han et al. (2012) J. Han, C. S. Frenk, V. R. Eke, L. Gao, S. D. M. White, A. Boyarsky, D. Malyshev, and O. Ruchayskiy, MNRAS 427, 1651 (2012), eprint 1207.6749.
- Hektor et al. (2013) A. Hektor, M. Raidal, and E. Tempel, Astrophys. J. Lett. 762, L22 (2013), eprint 1207.4466.
- Anderson et al. (2016) B. Anderson, S. Zimmer, J. Conrad, M. Gustafsson, M. Sánchez-Conde, and R. Caputo, JCAP 2016, 026 (2016), eprint 1511.00014.
- Liang et al. (2016) Y.-F. Liang, Z.-Q. Shen, X. Li, Y.-Z. Fan, X. Huang, S.-J. Lei, L. Feng, E.-W. Liang, and J. Chang, Phys. Rev. D 93, 103525 (2016), eprint 1602.06527.
- Ackermann et al. (2015) M. Ackermann, M. Ajello, A. Albert, W. B. Atwood, L. Baldini, G. Barbiellini, D. Bastieri, K. Bechtol, R. Bellazzini, E. Bissaldi, et al., Astrophys. J. 812, 159 (2015), eprint 1510.00004.
- Lisanti et al. (2018) M. Lisanti, S. Mishra-Sharma, N. L. Rodd, and B. R. Safdi, Phys. Rev. Lett. 120, 101101 (2018), eprint 1708.09385.
- Thorpe-Morgan et al. (2021) C. Thorpe-Morgan, D. Malyshev, C.-A. Stegen, A. Santangelo, and J. Jochum, MNRAS 502, 4039 (2021), eprint 2010.11006.
- Shen et al. (2021) Z.-Q. Shen, Z.-Q. Xia, and Y.-Z. Fan, Astrophys. J. 920, 1 (2021), eprint 2108.00363.
- Di Mauro et al. (2023) M. Di Mauro, J. Pérez-Romero, M. A. Sánchez-Conde, and N. Fornengo, Phys. Rev. D 107, 083030 (2023), eprint 2303.16930.
- Song et al. (2024) D. Song, N. Hiroshima, and K. Murase, arXiv e-prints arXiv:2401.15606 (2024), eprint 2401.15606.
- Tan et al. (2020) X. Tan, M. Colavincenzo, and S. Ammazzalorso, MNRAS 495, 114 (2020), eprint 1907.06905.
- Sunyaev and Zeldovich (1972) R. A. Sunyaev and Y. B. Zeldovich, Comments on Astrophysics and Space Physics 4, 173 (1972).
- Manna and Desai (2024a) S. Manna and S. Desai, JCAP 2024, 017 (2024a), eprint 2310.07519.
- Carlstrom et al. (2011) J. E. Carlstrom, P. A. R. Ade, K. A. Aird, B. A. Benson, L. E. Bleem, S. Busetti, C. L. Chang, E. Chauvin, H. M. Cho, T. M. Crawford, et al., Pub. Astro. Soc. Pac. 123, 568 (2011), eprint 0907.4445.
- Bleem et al. (2015a) L. E. Bleem, B. Stalder, T. de Haan, K. A. Aird, S. W. Allen, D. E. Applegate, M. L. N. Ashby, M. Bautz, M. Bayliss, B. A. Benson, et al., Astrophys. J. Suppl. Ser. 216, 27 (2015a), eprint 1409.0850.
- Bocquet et al. (2019) S. Bocquet, J. P. Dietrich, T. Schrabback, L. E. Bleem, M. Klein, S. W. Allen, D. E. Applegate, M. L. N. Ashby, M. Bautz, M. Bayliss, et al., Astrophys. J. 878, 55 (2019), eprint 1812.01679.
- Song et al. (2012) J. Song, A. Zenteno, B. Stalder, S. Desai, L. E. Bleem, K. A. Aird, R. Armstrong, M. L. N. Ashby, M. Bayliss, G. Bazin, et al., Astrophys. J. 761, 22 (2012), eprint 1207.4369.
- Ruel et al. (2014) J. Ruel, G. Bazin, M. Bayliss, M. Brodwin, R. J. Foley, B. Stalder, K. A. Aird, R. Armstrong, M. L. N. Ashby, M. Bautz, et al., Astrophys. J. 792, 45 (2014), eprint 1311.4953.
- Desai et al. (2012) S. Desai, R. Armstrong, J. J. Mohr, D. R. Semler, J. Liu, E. Bertin, S. S. Allam, W. A. Barkhouse, G. Bazin, E. J. Buckley-Geer, et al., Astrophys. J. 757, 83 (2012), eprint 1204.1210.
- Saro et al. (2015) A. Saro, S. Bocquet, E. Rozo, B. A. Benson, J. Mohr, E. S. Rykoff, M. Soares-Santos, L. Bleem, S. Dodelson, P. Melchior, et al., MNRAS 454, 2305 (2015), eprint 1506.07814.
- Charbonnier et al. (2012) A. Charbonnier, C. Combet, and D. Maurin, Computer Physics Communications 183, 656 (2012), eprint 1201.4728.
- Bonnivard et al. (2016) V. Bonnivard, M. Hütten, E. Nezri, A. Charbonnier, C. Combet, and D. Maurin, Computer Physics Communications 200, 336 (2016), eprint 1506.07628.
- Hütten et al. (2019) M. Hütten, C. Combet, and D. Maurin, Computer Physics Communications 235, 336 (2019), eprint 1806.08639.
- Cirelli et al. (2011) M. Cirelli, G. Corcella, A. Hektor, G. Hütsi, M. Kadastik, P. Panci, M. Raidal, F. Sala, and A. Strumia, JCAP 2011, 051 (2011), eprint 1012.4515.
- Pace and Strigari (2019) A. B. Pace and L. E. Strigari, MNRAS 482, 3480 (2019), eprint 1802.06811.
- Prada et al. (2012) F. Prada, A. A. Klypin, A. J. Cuesta, J. E. Betancort-Rijo, and J. Primack, MNRAS 423, 3018 (2012), eprint 1104.5130.
- Klypin et al. (2011) A. A. Klypin, S. Trujillo-Gomez, and J. Primack, Astrophys. J. 740, 102 (2011), eprint 1002.3660.
- Gao et al. (2004) L. Gao, S. D. M. White, A. Jenkins, F. Stoehr, and V. Springel, MNRAS 355, 819 (2004), eprint astro-ph/0404589.
- Kuhlen et al. (2012) M. Kuhlen, M. Vogelsberger, and R. Angulo, Physics of the Dark Universe 1, 50 (2012), eprint 1209.5745.
- Zavala and Frenk (2019) J. Zavala and C. S. Frenk, Galaxies 7, 81 (2019), eprint 1907.11775.
- Navarro et al. (1996) J. F. Navarro, C. S. Frenk, and S. D. M. White, Astrophys. J. 462, 563 (1996), eprint astro-ph/9508025.
- Navarro et al. (1997) J. F. Navarro, C. S. Frenk, and S. D. M. White, Astrophys. J. 490, 493 (1997), eprint astro-ph/9611107.
- Zhao (1996) H. Zhao, MNRAS 278, 488 (1996), eprint astro-ph/9509122.
- Bleem et al. (2015b) L. E. Bleem, B. Stalder, T. de Haan, K. A. Aird, S. W. Allen, D. E. Applegate, M. L. N. Ashby, M. Bautz, M. Bayliss, B. A. Benson, et al., Astrophys. J. Suppl. Ser. 216, 27 (2015b), eprint 1409.0850.
- White (2001) M. White, Astron. & Astrophys. 367, 27 (2001), eprint astro-ph/0011495.
- Sánchez-Conde and Prada (2014) M. A. Sánchez-Conde and F. Prada, MNRAS 442, 2271 (2014), eprint 1312.1729.
- Frenk and White (2012) C. S. Frenk and S. D. M. White, Annalen der Physik 524, 507 (2012), eprint 1210.0544.
- Hütten et al. (2018) M. Hütten, C. Combet, and D. Maurin, JCAP 2018, 005 (2018), eprint 1711.08323.
- Einasto (1965) J. Einasto, Trudy Astrofizicheskogo Instituta Alma-Ata 5, 87 (1965).
- Aguirre-Santaella et al. (2023) A. Aguirre-Santaella, M. A. Sánchez-Conde, G. Ogiya, J. Stücker, and R. E. Angulo, MNRAS 518, 93 (2023), eprint 2207.08652.
- Errani and Navarro (2021) R. Errani and J. F. Navarro, MNRAS 505, 18 (2021), eprint 2011.07077.
- Springel et al. (2008) V. Springel, J. Wang, M. Vogelsberger, A. Ludlow, A. Jenkins, A. Helmi, J. F. Navarro, C. S. Frenk, and S. D. M. White, MNRAS 391, 1685 (2008), eprint 0809.0898.
- Errani and Peñarrubia (2020) R. Errani and J. Peñarrubia, MNRAS 491, 4591 (2020), eprint 1906.01642.
- Kazantzidis et al. (2004) S. Kazantzidis, J. Magorrian, and B. Moore, Astrophys. J. 601, 37 (2004), eprint astro-ph/0309517.
- Peñarrubia et al. (2010) J. Peñarrubia, A. J. Benson, M. G. Walker, G. Gilmore, A. W. McConnachie, and L. Mayer, MNRAS 406, 1290 (2010), eprint 1002.3376.
- Moliné et al. (2017) Á. Moliné, M. A. Sánchez-Conde, S. Palomares-Ruiz, and F. Prada, MNRAS 466, 4974 (2017), eprint 1603.04057.
- Moliné et al. (2023) Á. Moliné, M. A. Sánchez-Conde, A. Aguirre-Santaella, T. Ishiyama, F. Prada, S. A. Cora, D. Croton, E. Jullo, R. B. Metcalf, T. Oogi, et al., MNRAS 518, 157 (2023), eprint 2110.02097.
- Newton et al. (2022) O. Newton, N. I. Libeskind, A. Knebe, M. A. Sánchez-Conde, J. G. Sorce, S. Pilipenko, M. Steinmetz, R. Pakmor, E. Tempel, Y. Hoffman, et al., MNRAS 514, 3612 (2022), eprint 2104.11242.
- Ackermann et al. (2012a) M. Ackermann, M. Ajello, A. Albert, A. Allafort, W. B. Atwood, M. Axelsson, L. Baldini, J. Ballet, G. Barbiellini, D. Bastieri, et al., Astrophys. J. Suppl. Ser. 203, 4 (2012a), eprint 1206.1896.
- Ackermann et al. (2012b) M. Ackermann, M. Ajello, W. B. Atwood, L. Baldini, J. Ballet, G. Barbiellini, D. Bastieri, K. Bechtol, R. Bellazzini, B. Berenji, et al., Astrophys. J. 750, 3 (2012b), eprint 1202.4039.
- Atwood et al. (2013) W. Atwood, A. Albert, L. Baldini, M. Tinivella, J. Bregeon, M. Pesce-Rollins, C. Sgrò, P. Bruel, E. Charles, A. Drlica-Wagner, et al., arXiv e-prints arXiv:1303.3514 (2013), eprint 1303.3514.
- Ballet et al. (2023) J. Ballet, P. Bruel, T. H. Burnett, B. Lott, and The Fermi-LAT collaboration, arXiv e-prints arXiv:2307.12546 (2023), eprint 2307.12546.
- Wood et al. (2017) M. Wood, R. Caputo, E. Charles, M. Di Mauro, J. Magill, J. S. Perkins, and Fermi-LAT Collaboration, in 35th International Cosmic Ray Conference (ICRC2017) (2017), vol. 301 of International Cosmic Ray Conference, p. 824, eprint 1707.09551.
- Jeltema and Profumo (2008) T. E. Jeltema and S. Profumo, JCAP 2008, 003 (2008), eprint 0808.2641.
- Mattox et al. (1996) J. R. Mattox, D. L. Bertsch, J. Chiang, B. L. Dingus, S. W. Digel, J. A. Esposito, J. M. Fierro, R. C. Hartman, S. D. Hunter, G. Kanbach, et al., Astrophys. J. 461, 396 (1996).
- Wilks (1938) S. S. Wilks, Annals Math. Statist. 9, 60 (1938).
- Pasumarti and Desai (2022) V. Pasumarti and S. Desai, JCAP 2022, 002 (2022), eprint 2210.12804.
- Manna and Desai (2024b) S. Manna and S. Desai, JCAP 2024, 013 (2024b), eprint 2401.13240.
- Di Mauro and Winkler (2021) M. Di Mauro and M. W. Winkler, Phys. Rev. D 103, 123005 (2021), eprint 2101.11027.
- Kluge et al. (2024) M. Kluge, J. Comparat, A. Liu, F. Balzer, E. Bulbul, J. Ider Chitham, V. Ghirardini, C. Garrel, Y. E. Bahar, E. Artis, et al., arXiv e-prints arXiv:2402.08453 (2024), eprint 2402.08453.
- Bleem et al. (2024) L. E. Bleem, M. Klein, T. M. C. Abbot, P. A. R. Ade, M. Aguena, O. Alves, A. J. Anderson, F. Andrade-Oliveira, B. Ansarinejad, M. Archipley, et al., The Open Journal of Astrophysics 7, 13 (2024), eprint 2311.07512.
Cluster Name | R.A. | Dec. | TS Values | |||||
(∘) | (∘) | (Kpc) | ||||||
SPT-CL J2012-5649 | 303.11 | -56.83 | 7.22 | 0.06 | 416.3 | 6.2 | ||
SPT-CL J0431-6126 | 67.84 | -61.44 | 7.2 | 0.06 | 416.0 | 1.5 | ||
SPT-CL J2313-4243 | 348.5 | -42.73 | 6.71 | 0.06 | 405.8 | 1.9 | ||
SPT-CL J2009-4518 | 302.45 | -45.31 | 6.23 | 0.06 | 395.2 | 1.9 | ||
SPT-CL J2201-5956 | 330.47 | -59.94 | 14.03 | 0.1 | 523.0 | 0.8 | ||
SPT-CL J0328-5541 | 52.17 | -55.7 | 8.18 | 0.08 | 435.0 | 3.4 | ||
SPT-CL J2217-6509 | 334.49 | -65.15 | 7.99 | 0.1 | 431.3 | 0.8 | ||
SPT-CL J2249-6426 | 342.43 | -64.43 | 7.22 | 0.09 | 416.3 | 1.5 | ||
SPT-CL J0145-5301 | 26.26 | -53.03 | 8.43 | 0.12 | 439.6 | 2.7 | ||
SPT-CL J0645-5413 | 101.37 | -54.22 | 15.33 | 0.16 | 538.8 | 1.5 | ||
SPT-CL J0500-5116 | 75.24 | -51.27 | 6.66 | 0.11 | 404.6 | 3.1 | ||
SPT-CL J2055-5456 | 314.0 | -54.94 | 8.48 | 0.14 | 440.5 | 2.4 | ||
SPT-CL J0628-4143 | 97.2 | -41.73 | 12.88 | 0.18 | 508.3 | 2.8 | ||
SPT-CL J0404-6510 | 61.05 | -65.18 | 5.74 | 0.12 | 383.9 | 2.5 | ||
SPT-CL J0641-5001 | 100.46 | -50.02 | 5.85 | 0.12 | 386.4 | 3.0 | ||
SPT-CL J0411-6340 | 62.86 | -63.68 | 7.29 | 0.14 | 417.7 | 2.5 | ||
SPT-CL J0145-6033 | 26.3 | -60.56 | 10.89 | 0.18 | 480.1 | 0.9 | ||
SPT-CL J2021-5257 | 305.48 | -52.95 | 6.36 | 0.14 | 398.1 | 9.0 | ||
SPT-CL J0638-5358 | 99.7 | -53.97 | 17.38 | 0.23 | 561.5 | 2.2 | ||
SPT-CL J2012-4130 | 303.0 | -41.5 | 7.01 | 0.15 | 412.1 | 1.2 | ||
SPT-CL J0516-6312 | 79.09 | -63.21 | 5.98 | 0.14 | 389.6 | 2.9 | ||
SPT-CL J2259-5617 | 345.0 | -56.29 | 7.09 | 0.15 | 413.6 | 1.9 | ||
SPT-CL J0027-5015 | 6.82 | -50.25 | 6.34 | 0.15 | 397.8 | 2.4 | ||
SPT-CL J0525-4715 | 81.46 | -47.26 | 11.0 | 0.19 | 481.7 | 2.7 | ||
SPT-CL J2254-5805 | 343.59 | -58.09 | 6.72 | 0.15 | 405.9 | 1.9 | ||
SPT-CL J0537-6504 | 84.35 | -65.07 | 10.67 | 0.2 | 476.9 | 2.7 | ||
SPT-CL J0658-5556 | 104.63 | -55.95 | 23.99 | 0.3 | 623.2 | 1.5 | ||
SPT-CL J0510-4519 | 77.58 | -45.33 | 10.55 | 0.2 | 474.9 | 0.8 | ||
SPT-CL J0637-4829 | 99.35 | -48.49 | 10.42 | 0.2 | 472.9 | 2.2 | ||
SPT-CL J2134-4238 | 323.5 | -42.64 | 9.44 | 0.2 | 457.1 | 2.4 | ||
SPT-CL J0051-4834 | 12.79 | -48.58 | 8.61 | 0.19 | 442.8 | 2.3 | ||
SPT-CL J0317-4849 | 49.45 | -48.83 | 6.08 | 0.16 | 391.9 | 1.5 | ||
SPT-CL J2023-5535 | 305.84 | -55.59 | 11.87 | 0.23 | 494.5 | 1.9 | ||
SPT-CL J0232-4421 | 38.07 | -44.35 | 17.49 | 0.28 | 562.6 | 2.2 | ||
SPT-CL J0216-4816 | 34.07 | -48.28 | 6.11 | 0.17 | 392.4 | 2.0 | ||
SPT-CL J2025-5117 | 306.48 | -51.29 | 10.05 | 0.22 | 467.1 | 2.1 | ||
SPT-CL J0010-5112 | 2.74 | -51.21 | 5.87 | 0.17 | 386.9 | 1.6 | ||
SPT-CL J2254-6314 | 343.51 | -63.25 | 9.02 | 0.21 | 450.1 | 1.4 | ||
SPT-CL J2248-4431 | 342.19 | -44.53 | 24.65 | 0.35 | 628.6 | 1.1 | ||
SPT-CL J0013-4621 | 3.47 | -46.36 | 5.88 | 0.18 | 387.1 | 2.8 | ||
SPT-CL J2020-4646 | 305.19 | -46.77 | 6.33 | 0.19 | 397.5 | 1.4 | ||
SPT-CL J0504-4929 | 76.01 | -49.49 | 6.82 | 0.2 | 408.2 | 1.2 | ||
SPT-CL J0225-4155 | 36.48 | -41.92 | 7.88 | 0.22 | 429.4 | 1.4 | ||
SPT-CL J0458-5741 | 74.6 | -57.7 | 5.82 | 0.19 | 385.8 | 1.9 | ||
SPT-CL J2241-4236 | 340.47 | -42.6 | 6.4 | 0.2 | 399.1 | 2.2 | ||
SPT-CL J0108-4341 | 17.13 | -43.69 | 5.74 | 0.19 | 383.9 | 1.8 | ||
SPT-CL J0124-5937 | 21.2 | -59.63 | 7.02 | 0.21 | 412.3 | 4.0 | ||
SPT-CL J0256-4736 | 44.24 | -47.61 | 8.02 | 0.23 | 432.0 | 2.1 | ||
SPT-CL J0653-5744 | 103.33 | -57.75 | 8.68 | 0.24 | 444.0 | 1.1 | ||
SPT-CL J2211-4833 | 332.83 | -48.56 | 8.33 | 0.24 | 437.7 | 2.3 | ||
SPT-CL J0118-5638 | 19.54 | -56.63 | 6.24 | 0.21 | 395.5 | 1.6 | ||
SPT-CL J0651-4037 | 102.82 | -40.63 | 7.8 | 0.24 | 427.9 | 3.5 | ||
SPT-CL J2121-6335 | 320.43 | -63.58 | 6.34 | 0.22 | 397.7 | 1.7 | ||
SPT-CL J0001-6258 | 0.4 | -62.98 | 5.85 | 0.21 | 386.5 | 1.1 | ||
SPT-CL J0235-5121 | 38.95 | -51.35 | 10.03 | 0.28 | 466.8 | 1.2 | ||
SPT-CL J2005-5635 | 301.34 | -56.59 | 5.58 | 0.21 | 380.0 | 2.5 | ||
SPT-CL J2254-4620 | 343.59 | -46.34 | 9.05 | 0.27 | 450.5 | 0.9 | ||
SPT-CL J0549-6205 | 87.33 | -62.09 | 18.2 | 0.38 | 570.1 | 1.1 | ||
SPT-CL J0516-5430 | 79.15 | -54.51 | 10.92 | 0.3 | 480.7 | 3.7 | ||
SPT-CL J2031-4037 | 307.97 | -40.62 | 14.64 | 0.34 | 530.5 | 1.8 | ||
SPT-CL J0225-4327 | 36.3 | -43.46 | 6.41 | 0.23 | 399.3 | 2.0 | ||
SPT-CL J0040-4407 | 10.2 | -44.13 | 15.23 | 0.35 | 537.5 | 1.5 | ||
SPT-CL J0620-4715 | 95.1 | -47.26 | 6.37 | 0.23 | 398.3 | 1.6 | ||
SPT-CL J2019-5642 | 304.77 | -56.71 | 6.11 | 0.23 | 392.5 | 3.0 | ||
SPT-CL J0249-5658 | 42.41 | -56.98 | 6.33 | 0.24 | 397.4 | 2.0 | ||
SPT-CL J0603-4714 | 90.99 | -47.24 | 8.58 | 0.27 | 442.4 | 2.3 | ||
SPT-CL J0601-4122 | 90.5 | -41.37 | 5.9 | 0.23 | 387.6 | 2.4 | ||
SPT-CL J2138-6008 | 324.51 | -60.13 | 11.18 | 0.32 | 484.5 | 0.8 | ||
SPT-CL J2032-5627 | 308.08 | -56.46 | 8.76 | 0.28 | 445.4 | 1.5 | ||
SPT-CL J2223-5015 | 335.81 | -50.27 | 6.13 | 0.24 | 392.9 | 2.6 | ||
SPT-CL J0412-5106 | 63.23 | -51.11 | 5.99 | 0.24 | 389.7 | 2.3 | ||
SPT-CL J2120-4016 | 320.14 | -40.27 | 6.35 | 0.25 | 397.9 | 2.6 | ||
SPT-CL J2219-5708 | 334.96 | -57.14 | 9.66 | 0.30 | 460.9 | 2.9 | ||
SPT-CL J2300-5331 | 345.18 | -53.52 | 6.78 | 0.26 | 407.2 | 4.6 | ||
SPT-CL J2027-4240 | 306.93 | -42.67 | 5.96 | 0.25 | 389.0 | 2.3 | ||
SPT-CL J0555-6406 | 88.87 | -64.1 | 11.79 | 0.35 | 493.3 | 1.3 | ||
SPT-CL J0311-6354 | 47.83 | -63.91 | 7.83 | 0.28 | 428.4 | 2.0 | ||
SPT-CL J0129-6432 | 22.43 | -64.54 | 10.34 | 0.33 | 471.7 | 2.1 | ||
SPT-CL J0505-6145 | 76.4 | -61.75 | 7.85 | 0.29 | 428.7 | 1.2 | ||
SPT-CL J0133-6434 | 23.41 | -64.57 | 9.59 | 0.32 | 459.7 | 1.3 | ||
SPT-CL J2325-4111 | 351.3 | -41.2 | 11.62 | 0.36 | 490.8 | 0.0 | ||
SPT-CL J0438-5419 | 69.57 | -54.32 | 16.08 | 0.42 | 547.3 | 2.6 | ||
SPT-CL J0405-4916 | 61.49 | -49.27 | 7.75 | 0.30 | 426.8 | 1.0 | ||
SPT-CL J2344-4224 | 356.15 | -42.41 | 5.65 | 0.26 | 381.7 | 2.0 | ||
SPT-CL J0150-4511 | 27.65 | -45.19 | 8.06 | 0.31 | 432.7 | 1.9 | ||
SPT-CL J0041-4428 | 10.25 | -44.48 | 9.29 | 0.33 | 454.6 | 0.7 | ||
SPT-CL J0214-4638 | 33.7 | -46.65 | 7.46 | 0.30 | 421.2 | 2.7 | ||
SPT-CL J0143-4452 | 25.89 | -44.87 | 6.06 | 0.27 | 391.3 | 2.4 | ||
SPT-CL J0655-5541 | 103.91 | -55.69 | 6.74 | 0.29 | 406.3 | 2.1 | ||
SPT-CL J2101-5542 | 315.31 | -55.7 | 5.77 | 0.27 | 384.5 | 2.4 | ||
SPT-CL J0440-4657 | 70.23 | -46.97 | 7.7 | 0.31 | 426.0 | 0.7 | ||
SPT-CL J0106-5943 | 16.62 | -59.72 | 9.85 | 0.35 | 463.9 | 1.5 | ||
SPT-CL J2130-6458 | 322.73 | -64.98 | 7.98 | 0.32 | 431.3 | 1.9 | ||
SPT-CL J0439-4600 | 69.81 | -46.01 | 8.55 | 0.33 | 441.8 | 2.7 | ||
SPT-CL J2011-5725 | 302.85 | -57.42 | 6.04 | 0.28 | 390.9 | 3.2 | ||
SPT-CL J0348-4515 | 57.07 | -45.25 | 9.75 | 0.36 | 462.2 | 2.4 | ||
SPT-CL J0304-4921 | 46.06 | -49.36 | 11.67 | 0.39 | 491.6 | 2.0 | ||
SPT-CL J0455-4159 | 74.0 | -41.99 | 6.07 | 0.29 | 391.6 | 3.5 | ||
SPT-CL J2223-5227 | 335.87 | -52.47 | 6.07 | 0.29 | 391.5 | 0.8 | ||
SPT-CL J2115-4659 | 318.8 | -46.99 | 6.46 | 0.30 | 400.2 | 1.3 | ||
SPT-CL J2016-4954 | 304.01 | -49.91 | 5.93 | 0.29 | 388.3 | 1.8 | ||
SPT-CL J0522-4818 | 80.57 | -48.3 | 6.2 | 0.30 | 394.5 | 3.1 | ||
SPT-CL J0151-5654 | 27.79 | -56.91 | 5.84 | 0.29 | 386.2 | 2.7 | ||
SPT-CL J0234-5831 | 38.68 | -58.52 | 12.32 | 0.42 | 500.7 | 2.4 | ||
SPT-CL J0114-4123 | 18.68 | -41.4 | 10.87 | 0.39 | 479.9 | 2.7 | ||
SPT-CL J0022-4144 | 5.55 | -41.74 | 6.14 | 0.30 | 393.1 | 3.5 | ||
SPT-CL J0411-4819 | 62.82 | -48.32 | 12.51 | 0.42 | 503.3 | 2.1 | ||
SPT-CL J0551-4339 | 87.88 | -43.66 | 7.36 | 0.33 | 419.2 | 3.1 | ||
SPT-CL J2355-5055 | 358.95 | -50.93 | 6.84 | 0.32 | 408.5 | 3.0 | ||
SPT-CL J0509-6118 | 77.47 | -61.31 | 10.27 | 0.39 | 470.6 | 2.2 | ||
SPT-CL J0636-4942 | 99.17 | -49.7 | 8.27 | 0.35 | 436.7 | 3.5 | ||
SPT-CL J0110-4445 | 17.59 | -44.76 | 8.55 | 0.36 | 441.8 | 2.2 | ||
SPT-CL J0013-4906 | 3.33 | -49.12 | 11.04 | 0.41 | 482.4 | 2.5 | ||
SPT-CL J0001-4842 | 0.28 | -48.71 | 6.83 | 0.33 | 408.2 | 1.7 | ||
SPT-CL J0217-5245 | 34.3 | -52.76 | 7.36 | 0.34 | 419.3 | 1.3 | ||
SPT-CL J0254-5857 | 43.57 | -58.95 | 12.01 | 0.44 | 496.5 | 2.7 | ||
SPT-CL J2330-4502 | 352.57 | -45.03 | 6.17 | 0.32 | 393.8 | 2.3 | ||
SPT-CL J0304-4401 | 46.07 | -44.03 | 13.07 | 0.46 | 510.8 | 2.4 | ||
SPT-CL J0236-4938 | 39.25 | -49.64 | 6.74 | 0.33 | 406.3 | 1.1 | ||
SPT-CL J0600-4353 | 90.06 | -43.89 | 7.79 | 0.36 | 427.7 | 2.0 | ||
SPT-CL J0052-5657 | 13.16 | -56.96 | 5.95 | 0.32 | 388.9 | 0.7 | ||
SPT-CL J0650-4503 | 102.68 | -45.06 | 9.61 | 0.40 | 459.9 | 2.1 | ||
SPT-CL J2059-5018 | 314.93 | -50.31 | 6.16 | 0.33 | 393.7 | 2.5 | ||
SPT-CL J2131-4019 | 322.77 | -40.32 | 11.66 | 0.45 | 491.4 | 0.6 | ||
SPT-CL J0416-6359 | 64.16 | -64.0 | 6.83 | 0.35 | 408.2 | 2.0 | ||
SPT-CL J0424-4406 | 66.0 | -44.11 | 7.24 | 0.36 | 416.8 | 1.3 | ||
SPT-CL J0240-5946 | 40.16 | -59.77 | 8.92 | 0.4 | 448.3 | 2.5 | ||
SPT-CL J0144-4807 | 26.18 | -48.13 | 5.84 | 0.33 | 386.2 | 0.9 | ||
SPT-CL J0405-4648 | 61.29 | -46.81 | 7.2 | 0.36 | 416.0 | 1.1 | ||
SPT-CL J2358-6129 | 359.71 | -61.49 | 6.79 | 0.36 | 407.5 | 2.0 | ||
SPT-CL J2135-5726 | 323.92 | -57.44 | 9.77 | 0.43 | 462.7 | 3.2 | ||
SPT-CL J0330-5228 | 52.73 | -52.47 | 10.48 | 0.44 | 473.8 | 1.2 | ||
SPT-CL J0412-5743 | 63.02 | -57.72 | 6.06 | 0.34 | 391.4 | 1.6 | ||
SPT-CL J2332-5358 | 353.11 | -53.97 | 8.54 | 0.40 | 441.6 | 1.4 | ||
SPT-CL J0012-5352 | 3.06 | -53.87 | 5.95 | 0.34 | 388.8 | 1.9 | ||
SPT-CL J2205-5927 | 331.27 | -59.46 | 6.66 | 0.37 | 404.6 | 0.5 | ||
SPT-CL J2344-4243 | 356.18 | -42.72 | 18.1 | 0.6 | 569.1 | 1.6 | ||
SPT-CL J0402-4611 | 60.58 | -46.19 | 6.2 | 0.36 | 394.6 | 1.8 | ||
SPT-CL J0052-4551 | 13.19 | -45.86 | 6.2 | 0.36 | 394.4 | 2.0 | ||
SPT-CL J0243-4833 | 40.91 | -48.56 | 12.1 | 0.5 | 497.7 | 1.0 | ||
SPT-CL J2022-6323 | 305.53 | -63.4 | 6.99 | 0.38 | 411.6 | 3.0 | ||
SPT-CL J2327-5137 | 351.78 | -51.62 | 5.39 | 0.34 | 375.2 | 2.4 | ||
SPT-CL J2316-5453 | 349.21 | -54.9 | 6.48 | 0.37 | 400.8 | 1.6 | ||
SPT-CL J0344-5518 | 56.21 | -55.3 | 5.94 | 0.36 | 388.5 | 1.8 | ||
SPT-CL J2145-5644 | 326.47 | -56.75 | 10.87 | 0.48 | 479.8 | 2.6 | ||
SPT-CL J2233-5339 | 338.33 | -53.65 | 9.0 | 0.44 | 449.6 | 1.0 | ||
SPT-CL J2159-6244 | 329.99 | -62.74 | 6.96 | 0.39 | 411.0 | 1.6 | ||
SPT-CL J0445-4230 | 71.28 | -42.51 | 7.77 | 0.41 | 427.2 | 1.9 | ||
SPT-CL J0551-5709 | 87.9 | -57.16 | 8.13 | 0.42 | 434.1 | 0.7 | ||
SPT-CL J2124-6124 | 321.15 | -61.41 | 8.45 | 0.44 | 439.9 | 2.7 | ||
SPT-CL J2351-5452 | 357.9 | -54.88 | 6.5 | 0.38 | 401.2 | 0.5 | ||
SPT-CL J2206-4057 | 331.62 | -40.95 | 5.63 | 0.36 | 381.1 | 1.1 | ||
SPT-CL J0252-4824 | 43.19 | -48.41 | 7.76 | 0.42 | 427.0 | 0.0 | ||
SPT-CL J0429-4355 | 67.31 | -43.93 | 5.85 | 0.37 | 386.3 | 0.8 | ||
SPT-CL J2259-5431 | 344.98 | -54.53 | 6.47 | 0.39 | 400.5 | 1.2 | ||
SPT-CL J0025-5034 | 6.37 | -50.57 | 5.88 | 0.38 | 387.0 | 0.3 | ||
SPT-CL J0447-5055 | 71.84 | -50.92 | 6.57 | 0.4 | 402.7 | 0.2 | ||
SPT-CL J0019-4051 | 4.76 | -40.86 | 9.61 | 0.48 | 459.9 | 1.0 | ||
SPT-CL J2136-4704 | 324.12 | -47.08 | 7.35 | 0.43 | 419.1 | 0.5 | ||
SPT-CL J0354-5904 | 58.56 | -59.07 | 7.02 | 0.42 | 412.2 | 1.4 | ||
SPT-CL J0333-5842 | 53.32 | -58.7 | 5.47 | 0.37 | 377.3 | 0.4 | ||
SPT-CL J0244-4857 | 41.03 | -48.96 | 6.76 | 0.41 | 406.9 | 0.3 | ||
SPT-CL J0342-4028 | 55.56 | -40.48 | 7.59 | 0.44 | 423.7 | 1.8 | ||
SPT-CL J2030-5638 | 307.7 | -56.64 | 6.07 | 0.39 | 391.5 | 0.5 | ||
SPT-CL J0509-5342 | 77.34 | -53.71 | 8.27 | 0.46 | 436.6 | 0.2 | ||
SPT-CL J2319-4716 | 349.98 | -47.28 | 6.98 | 0.43 | 411.5 | 1.7 | ||
SPT-CL J0611-5938 | 92.81 | -59.64 | 5.75 | 0.39 | 384.1 | 0.3 | ||
SPT-CL J0054-4046 | 13.59 | -40.78 | 6.3 | 0.41 | 396.7 | 1.3 | ||
SPT-CL J0505-4204 | 76.37 | -42.08 | 5.94 | 0.4 | 388.6 | 2.0 | ||
SPT-CL J0655-5234 | 103.96 | -52.57 | 8.39 | 0.47 | 438.9 | 1.0 | ||
SPT-CL J0334-4659 | 53.55 | -46.99 | 8.95 | 0.49 | 448.5 | 1.6 | ||
SPT-CL J2016-4517 | 304.0 | -45.3 | 5.9 | 0.4 | 387.6 | 1.1 | ||
SPT-CL J2342-4714 | 355.75 | -47.24 | 5.82 | 0.4 | 385.6 | 1.5 | ||
SPT-CL J0047-4506 | 11.82 | -45.11 | 8.11 | 0.47 | 433.7 | 1.3 | ||
SPT-CL J0439-5330 | 69.93 | -53.5 | 6.22 | 0.41 | 394.9 | 0.0 | ||
SPT-CL J0647-5828 | 101.98 | -58.48 | 7.14 | 0.44 | 414.7 | 1.7 | ||
SPT-CL J0341-5027 | 55.28 | -50.46 | 5.48 | 0.39 | 377.3 | 0.5 | ||
SPT-CL J0254-6051 | 43.6 | -60.86 | 7.09 | 0.44 | 413.7 | 0.9 | ||
SPT-CL J2140-5727 | 325.14 | -57.46 | 5.92 | 0.40 | 387.9 | 0.1 | ||
SPT-CL J0452-4806 | 73.0 | -48.11 | 5.29 | 0.38 | 372.7 | 0.5 | ||
SPT-CL J0259-4556 | 44.9 | -45.94 | 6.58 | 0.43 | 402.8 | 1.1 | ||
SPT-CL J0626-4446 | 96.74 | -44.77 | 6.92 | 0.44 | 410.2 | 0.2 | ||
SPT-CL J0237-4151 | 39.42 | -41.86 | 5.78 | 0.41 | 384.7 | 0.0 | ||
SPT-CL J2035-5251 | 308.8 | -52.85 | 9.97 | 0.53 | 465.8 | 1.5 | ||
SPT-CL J0014-4036 | 3.74 | -40.6 | 9.57 | 0.52 | 459.3 | 0.2 | ||
SPT-CL J0257-5842 | 44.39 | -58.71 | 6.0 | 0.42 | 389.8 | 0.0 | ||
SPT-CL J0638-4243 | 99.57 | -42.72 | 5.67 | 0.41 | 382.0 | 0.2 | ||
SPT-CL J0417-4748 | 64.35 | -47.81 | 11.66 | 0.58 | 491.3 | 1.6 | ||
SPT-CL J0655-4429 | 103.76 | -44.48 | 5.69 | 0.41 | 382.5 | 2.0 | ||
SPT-CL J0543-6219 | 85.76 | -62.33 | 8.59 | 0.51 | 442.4 | 1.0 | ||
SPT-CL J2335-4544 | 353.79 | -45.74 | 9.98 | 0.55 | 466.0 | 1.2 | ||
SPT-CL J2306-6505 | 346.73 | -65.09 | 9.29 | 0.53 | 454.7 | 0.0 | ||
SPT-CL J0611-4724 | 92.92 | -47.41 | 6.5 | 0.45 | 401.3 | 0.0 | ||
SPT-CL J2112-4434 | 318.21 | -44.58 | 9.13 | 0.53 | 451.9 | 0.7 | ||
SPT-CL J0038-5244 | 9.72 | -52.74 | 5.87 | 0.43 | 386.7 | 0.0 | ||
SPT-CL J2111-5339 | 317.92 | -53.65 | 6.45 | 0.45 | 400.1 | 0.0 | ||
SPT-CL J0351-5636 | 57.93 | -56.61 | 5.32 | 0.41 | 373.4 | 0.4 | ||
SPT-CL J0546-4752 | 86.55 | -47.88 | 6.04 | 0.43 | 390.9 | 0.7 | ||
SPT-CL J2050-4213 | 312.57 | -42.22 | 7.96 | 0.5 | 430.9 | 1.0 | ||
SPT-CL J0200-4852 | 30.14 | -48.88 | 7.93 | 0.5 | 430.2 | 0.0 | ||
SPT-CL J0216-4830 | 34.07 | -48.51 | 6.97 | 0.47 | 411.3 | 1.2 | ||
SPT-CL J0346-5439 | 56.72 | -54.65 | 8.91 | 0.53 | 448.1 | 0.0 | ||
SPT-CL J2259-3952 | 344.81 | -39.87 | 9.1 | 0.54 | 451.4 | 0.0 | ||
SPT-CL J0257-4817 | 44.45 | -48.3 | 5.96 | 0.44 | 389.0 | 0.0 | ||
SPT-CL J0451-4952 | 72.97 | -49.88 | 5.56 | 0.42 | 379.3 | 0.0 | ||
SPT-CL J0532-5450 | 83.03 | -54.84 | 5.69 | 0.43 | 382.7 | 0.0 | ||
SPT-CL J0317-5935 | 49.32 | -59.59 | 6.8 | 0.47 | 407.7 | 0.0 | ||
SPT-CL J0257-5732 | 44.35 | -57.54 | 5.72 | 0.43 | 383.4 | 0.0 | ||
SPT-CL J2235-4416 | 338.86 | -44.27 | 6.19 | 0.45 | 394.4 | 0.1 | ||
SPT-CL J0517-6311 | 79.41 | -63.2 | 5.99 | 0.45 | 389.6 | 0.2 | ||
SPT-CL J0403-5719 | 60.97 | -57.32 | 6.44 | 0.47 | 399.9 | 0.1 | ||
SPT-CL J0456-6141 | 74.15 | -61.68 | 5.51 | 0.44 | 378.3 | 0.0 | ||
SPT-CL J0314-6130 | 48.61 | -61.51 | 5.37 | 0.43 | 374.8 | 0.0 | ||
SPT-CL J0508-6149 | 77.16 | -61.82 | 5.73 | 0.45 | 383.5 | 0.3 | ||
SPT-CL J0641-5950 | 100.38 | -59.85 | 7.79 | 0.52 | 427.5 | 0.1 | ||
SPT-CL J0417-4427 | 64.41 | -44.46 | 9.08 | 0.56 | 450.9 | 0.1 | ||
SPT-CL J0307-5042 | 46.95 | -50.71 | 8.67 | 0.55 | 443.8 | 0.0 | ||
SPT-CL J0232-5257 | 38.19 | -52.96 | 8.79 | 0.56 | 446.1 | 0.4 | ||
SPT-CL J0124-4301 | 21.15 | -43.02 | 6.04 | 0.47 | 390.8 | 0.1 | ||
SPT-CL J2132-4349 | 323.17 | -43.83 | 7.48 | 0.52 | 421.6 | 0.4 | ||
SPT-CL J2131-5003 | 322.97 | -50.06 | 5.86 | 0.46 | 386.5 | 0.0 | ||
SPT-CL J0102-4915 | 15.73 | -49.26 | 22.05 | 0.87 | 606.6 | 1.9 | ||
SPT-CL J0113-6105 | 18.4 | -61.09 | 5.57 | 0.45 | 379.6 | 0.0 | ||
SPT-CL J0304-4748 | 46.15 | -47.81 | 7.06 | 0.51 | 413.1 | 0.0 | ||
SPT-CL J2331-5051 | 352.96 | -50.86 | 9.09 | 0.58 | 451.2 | 0.0 | ||
SPT-CL J0544-3950 | 86.25 | -39.85 | 7.42 | 0.52 | 420.3 | 0.0 | ||
SPT-CL J0048-4548 | 12.25 | -45.8 | 5.83 | 0.47 | 385.8 | 0.0 | ||
SPT-CL J0007-4706 | 1.75 | -47.12 | 5.59 | 0.46 | 380.2 | 0.2 | ||
SPT-CL J0306-4749 | 46.74 | -47.82 | 5.58 | 0.46 | 380.0 | 0.1 | ||
SPT-CL J0337-6300 | 54.47 | -63.01 | 5.78 | 0.47 | 384.8 | 0.0 | ||
SPT-CL J0111-5424 | 17.77 | -54.42 | 6.25 | 0.48 | 395.7 | 0.1 | ||
SPT-CL J0456-5116 | 74.12 | -51.28 | 8.4 | 0.56 | 439.0 | 0.0 | ||
SPT-CL J0257-6050 | 44.34 | -60.85 | 5.43 | 0.46 | 376.3 | 0.1 | ||
SPT-CL J0659-5300 | 104.78 | -53.01 | 6.0 | 0.48 | 389.9 | 0.0 | ||
SPT-CL J2145-4348 | 326.36 | -43.8 | 6.22 | 0.49 | 394.9 | 0.0 | ||
SPT-CL J0253-6046 | 43.46 | -60.77 | 5.59 | 0.47 | 380.2 | 0.0 | ||
SPT-CL J2302-4435 | 345.59 | -44.58 | 5.54 | 0.47 | 378.9 | 0.0 | ||
SPT-CL J2245-6206 | 341.26 | -62.11 | 8.87 | 0.59 | 447.4 | 0.0 | ||
SPT-CL J0337-4928 | 54.46 | -49.47 | 6.08 | 0.49 | 391.7 | 0.0 | ||
SPT-CL J2232-5959 | 338.15 | -59.99 | 9.1 | 0.59 | 451.4 | 0.2 | ||
SPT-CL J2040-5342 | 310.22 | -53.71 | 6.88 | 0.52 | 409.4 | 0.1 | ||
SPT-CL J0559-5249 | 89.93 | -52.83 | 9.37 | 0.61 | 455.9 | 0.0 | ||
SPT-CL J0307-6225 | 46.83 | -62.43 | 8.37 | 0.58 | 438.4 | 0.0 | ||
SPT-CL J0025-4133 | 6.49 | -41.55 | 7.2 | 0.54 | 415.9 | 0.0 | ||
SPT-CL J0109-4045 | 17.48 | -40.77 | 5.55 | 0.48 | 379.2 | 0.0 | ||
SPT-CL J0238-4904 | 39.7 | -49.07 | 6.46 | 0.52 | 400.4 | 0.0 | ||
SPT-CL J0221-4446 | 35.42 | -44.78 | 5.94 | 0.5 | 388.5 | 0.2 | ||
SPT-CL J2051-6256 | 312.8 | -62.93 | 5.54 | 0.48 | 378.9 | 0.1 | ||
SPT-CL J0218-4315 | 34.58 | -43.26 | 9.43 | 0.63 | 456.9 | 0.1 | ||
SPT-CL J0106-5355 | 16.57 | -53.93 | 5.82 | 0.5 | 385.7 | 0.0 | ||
SPT-CL J0219-4934 | 34.81 | -49.58 | 7.01 | 0.55 | 412.0 | 0.0 | ||
SPT-CL J0111-5518 | 17.84 | -55.31 | 5.44 | 0.49 | 376.6 | 0.0 | ||
SPT-CL J0351-4109 | 57.75 | -41.16 | 8.92 | 0.62 | 448.3 | 0.1 | ||
SPT-CL J2017-6258 | 304.48 | -62.98 | 6.62 | 0.54 | 403.8 | 0.3 | ||
SPT-CL J2148-6116 | 327.18 | -61.28 | 7.5 | 0.57 | 422.0 | 0.0 | ||
SPT-CL J0619-5802 | 94.92 | -58.04 | 7.0 | 0.55 | 411.9 | 0.2 | ||
SPT-CL J0135-5902 | 23.79 | -59.04 | 5.85 | 0.51 | 386.4 | 0.1 | ||
SPT-CL J0041-5107 | 10.29 | -51.13 | 5.57 | 0.5 | 379.8 | 0.0 | ||
SPT-CL J0342-5354 | 55.52 | -53.91 | 6.02 | 0.52 | 390.3 | 0.0 | ||
SPT-CL J0135-5904 | 23.98 | -59.08 | 5.41 | 0.5 | 375.8 | 0.1 | ||
SPT-CL J0343-5518 | 55.76 | -55.3 | 6.48 | 0.54 | 400.8 | 0.0 | ||
SPT-CL J0033-6326 | 8.48 | -63.45 | 7.93 | 0.6 | 430.3 | 0.1 | ||
SPT-CL J0212-4657 | 33.11 | -46.95 | 9.55 | 0.65 | 459.0 | 0.2 | ||
SPT-CL J0410-6343 | 62.52 | -63.73 | 6.11 | 0.53 | 392.5 | 0.0 | ||
SPT-CL J0336-4005 | 54.16 | -40.1 | 5.8 | 0.52 | 385.2 | 0.2 | ||
SPT-CL J0217-4310 | 34.41 | -43.18 | 7.07 | 0.57 | 413.3 | 0.0 | ||
SPT-CL J2218-4519 | 334.75 | -45.32 | 8.8 | 0.64 | 446.1 | 0.1 | ||
SPT-CL J0429-5233 | 67.43 | -52.56 | 5.21 | 0.5 | 370.6 | 0.0 | ||
SPT-CL J2337-5942 | 354.35 | -59.7 | 13.26 | 0.78 | 513.2 | 0.3 | ||
SPT-CL J2155-6048 | 328.99 | -60.81 | 6.11 | 0.54 | 392.3 | 0.0 | ||
SPT-CL J2350-5301 | 357.73 | -53.02 | 6.11 | 0.54 | 392.4 | 0.0 | ||
SPT-CL J0542-4100 | 85.72 | -41.0 | 8.61 | 0.64 | 442.8 | 0.0 | ||
SPT-CL J2222-4834 | 335.71 | -48.57 | 8.95 | 0.65 | 448.8 | 0.0 | ||
SPT-CL J0522-5026 | 80.52 | -50.44 | 5.55 | 0.52 | 379.2 | 0.3 | ||
SPT-CL J0309-4958 | 47.26 | -49.97 | 6.32 | 0.55 | 397.3 | 0.0 | ||
SPT-CL J0202-5401 | 30.58 | -54.02 | 6.31 | 0.55 | 397.1 | 0.0 | ||
SPT-CL J0426-5455 | 66.52 | -54.92 | 8.53 | 0.64 | 441.4 | 0.1 | ||
SPT-CL J0142-5032 | 25.55 | -50.54 | 9.55 | 0.68 | 458.9 | 0.4 | ||
SPT-CL J0402-6130 | 60.71 | -61.5 | 5.41 | 0.52 | 375.8 | 0.0 | ||
SPT-CL J0011-4614 | 2.98 | -46.24 | 6.0 | 0.54 | 389.8 | 0.3 | ||
SPT-CL J2155-5224 | 328.89 | -52.41 | 6.11 | 0.55 | 392.5 | 0.2 | ||
SPT-CL J0444-4352 | 71.17 | -43.87 | 5.7 | 0.53 | 383.0 | 0.1 | ||
SPT-CL J2020-6314 | 305.03 | -63.24 | 5.78 | 0.54 | 384.8 | 0.1 | ||
SPT-CL J0030-5213 | 7.53 | -52.22 | 5.58 | 0.53 | 379.9 | 0.1 | ||
SPT-CL J0512-5139 | 78.16 | -51.66 | 6.38 | 0.57 | 398.6 | 0.3 | ||
SPT-CL J0152-5303 | 28.23 | -53.05 | 7.34 | 0.61 | 418.8 | 0.0 | ||
SPT-CL J2206-5807 | 331.66 | -58.13 | 7.02 | 0.6 | 412.3 | 0.0 | ||
SPT-CL J0157-4007 | 29.45 | -40.13 | 6.4 | 0.57 | 398.9 | 0.1 | ||
SPT-CL J0649-4510 | 102.45 | -45.17 | 5.96 | 0.55 | 389.0 | 0.2 | ||
SPT-CL J0543-4250 | 85.94 | -42.84 | 7.58 | 0.62 | 423.6 | 0.0 | ||
SPT-CL J2110-5244 | 317.55 | -52.75 | 7.12 | 0.61 | 414.4 | 0.0 | ||
SPT-CL J0243-5930 | 40.86 | -59.51 | 7.7 | 0.64 | 426.0 | 0.0 | ||
SPT-CL J0217-5014 | 34.27 | -50.24 | 5.43 | 0.54 | 376.2 | 0.0 | ||
SPT-CL J0256-5617 | 44.1 | -56.3 | 7.55 | 0.64 | 423.0 | 0.0 | ||
SPT-CL J2312-4621 | 348.06 | -46.35 | 7.31 | 0.63 | 418.3 | 0.0 | ||
SPT-CL J2354-5633 | 358.71 | -56.55 | 5.78 | 0.56 | 384.9 | 0.0 | ||
SPT-CL J2220-4534 | 335.08 | -45.58 | 7.47 | 0.64 | 421.4 | 0.2 | ||
SPT-CL J0519-4248 | 79.85 | -42.81 | 5.75 | 0.57 | 383.9 | 0.0 | ||
SPT-CL J0422-5140 | 65.59 | -51.68 | 6.22 | 0.59 | 395.1 | 0.1 | ||
SPT-CL J2140-5331 | 325.03 | -53.52 | 5.67 | 0.57 | 382.2 | 0.1 | ||
SPT-CL J2146-5736 | 326.7 | -57.61 | 6.42 | 0.60 | 399.5 | 0.4 | ||
SPT-CL J0334-4815 | 53.71 | -48.26 | 7.3 | 0.64 | 417.9 | 0.2 | ||
SPT-CL J0148-4518 | 27.1 | -45.3 | 5.78 | 0.57 | 384.8 | 0.1 | ||
SPT-CL J2146-4846 | 326.53 | -48.78 | 6.79 | 0.62 | 407.5 | 0.3 | ||
SPT-CL J0123-4821 | 20.79 | -48.36 | 7.53 | 0.66 | 422.5 | 0.1 | ||
SPT-CL J0145-4426 | 26.28 | -44.44 | 5.53 | 0.57 | 378.6 | 0.2 | ||
SPT-CL J2317-4707 | 349.27 | -47.12 | 5.49 | 0.57 | 377.9 | 0.0 | ||
SPT-CL J0352-5647 | 58.24 | -56.8 | 7.28 | 0.65 | 417.6 | 0.1 | ||
SPT-CL J0231-5403 | 37.78 | -54.06 | 5.85 | 0.59 | 386.4 | 0.0 | ||
SPT-CL J0231-4427 | 37.75 | -44.46 | 6.0 | 0.6 | 389.9 | 0.0 | ||
SPT-CL J2109-5040 | 317.38 | -50.68 | 6.39 | 0.62 | 398.7 | 0.0 | ||
SPT-CL J0615-5746 | 93.96 | -57.78 | 5.82 | 0.97 | 552.8 | 0.0 | ||
SPT-CL J0511-5154 | 77.92 | -51.9 | 6.98 | 0.65 | 411.5 | 0.1 | ||
SPT-CL J0307-4123 | 46.88 | -41.39 | 5.82 | 0.59 | 385.6 | 0.1 | ||
SPT-CL J0216-4219 | 34.06 | -42.33 | 6.34 | 0.62 | 397.7 | 0.0 | ||
SPT-CL J0612-4317 | 93.02 | -43.3 | 6.74 | 0.64 | 406.4 | 0.0 | ||
SPT-CL J2118-5055 | 319.73 | -50.93 | 6.37 | 0.63 | 398.3 | 0.0 | ||
SPT-CL J0622-4645 | 95.71 | -46.77 | 5.88 | 0.60 | 387.1 | 0.0 | ||
SPT-CL J0341-6143 | 55.35 | -61.72 | 6.07 | 0.61 | 391.6 | 0.0 | ||
SPT-CL J0049-5315 | 12.38 | -53.25 | 7.01 | 0.66 | 412.1 | 0.0 | ||
SPT-CL J0000-5748 | 0.25 | -57.81 | 7.68 | 0.70 | 425.5 | 0.0 | ||
SPT-CL J0014-4952 | 3.7 | -49.88 | 8.86 | 0.75 | 447.3 | 0.0 | ||
SPT-CL J0521-5104 | 80.3 | -51.08 | 7.04 | 0.68 | 412.7 | 0.0 | ||
SPT-CL J0233-5819 | 38.26 | -58.33 | 6.78 | 0.66 | 407.2 | 0.0 | ||
SPT-CL J2134-4109 | 323.53 | -41.16 | 7.3 | 0.7 | 418.0 | 0.0 | ||
SPT-CL J0359-5218 | 59.84 | -52.32 | 6.49 | 0.66 | 401.0 | 0.0 | ||
SPT-CL J0528-4417 | 82.06 | -44.29 | 6.01 | 0.64 | 370.2 | 0.3 | ||
SPT-CL J0021-4902 | 5.38 | -49.04 | 6.04 | 0.64 | 390.8 | 0.0 | ||
SPT-CL J0310-4647 | 47.63 | -46.78 | 7.4 | 0.71 | 419.9 | 0.0 | ||
SPT-CL J2043-5035 | 310.83 | -50.59 | 7.74 | 0.72 | 426.7 | 0.0 | ||
SPT-CL J0324-6236 | 51.05 | -62.6 | 8.29 | 0.75 | 437.1 | 0.0 | ||
SPT-CL J0441-4855 | 70.45 | -48.92 | 8.09 | 0.74 | 433.3 | 0.0 | ||
SPT-CL J0406-4805 | 61.73 | -48.09 | 7.81 | 0.74 | 428.0 | 0.0 | ||
SPT-CL J2358-4354 | 359.73 | -43.90 | 6.29 | 0.67 | 396.6 | 0.0 | ||
SPT-CL J0230-6028 | 37.64 | -60.47 | 6.29 | 0.67 | 396.6 | 0.0 | ||
SPT-CL J2259-6057 | 344.75 | -60.95 | 9.16 | 0.81 | 452.4 | 0.0 | ||
SPT-CL J2034-5936 | 308.54 | -59.6 | 7.96 | 0.76 | 430.8 | 0.0 | ||
SPT-CL J0103-4250 | 15.91 | -42.83 | 6.95 | 0.71 | 410.7 | 0.0 | ||
SPT-CL J0118-5156 | 19.60 | -51.94 | 6.58 | 0.70 | 402.9 | 0.1 | ||
SPT-CL J0449-4901 | 72.27 | -49.02 | 8.25 | 0.79 | 436.3 | 0.0 | ||
SPT-CL J2106-4421 | 316.57 | -44.35 | 7.81 | 0.78 | 427.9 | 0.0 | ||
SPT-CL J0131-5604 | 22.93 | -56.08 | 6.95 | 0.75 | 373.1 | 0.1 | ||
SPT-CL J0641-4733 | 100.29 | -47.57 | 7.40 | 0.78 | 419.9 | 0.0 | ||
SPT-CL J2258–4044 | 344.71 | -40.74 | 9.65 | 0.90 | 460.6 | 0.0 | ||
SPT-CL J2301-4023 | 345.47 | -40.39 | 7.93 | 0.83 | 430.4 | 0.0 | ||
SPT-CL J0058-6145 | 14.58 | -61.76 | 7.59 | 0.69 | 423.7 | 0.1 | ||
SPT-CL J2146-4633 | 326.65 | -46.55 | 9.17 | 0.93 | 452.6 | 0.0 | ||
SPT-CL J0646-6236 | 101.64 | -62.61 | 8.26 | 0.89 | 436.5 | 0.0 | ||
SPT-CL J2106-5844 | 316.52 | -58.75 | 13.57 | 1.13 | 517.3 | 0.0 | ||
SPT-CL J2341-5119 | 355.30 | -51.33 | 9.36 | 1.00 | 455.8 | 0.0 |